lexxii:

show how the alternative definition of power p=fv can be derived by substituting the definitions of work and speed into the standard definition of power p=w/delta t

sillybilly123:

By definition of work: \(\delta W = \mathbf F \cdot \delta \mathbf r \) And so instantaneous Power at time \(t\) is: \(P = \lim\limits_{\delta t \to 0} \dfrac{\delta W}{\delta t} = \lim\limits_{\delta t \to 0} \dfrac{\mathbf F \cdot \delta \mathbf r}{\delta t} \) \( = \mathbf F \cdot \lim\limits_{\delta t \to 0} \dfrac{ \delta \mathbf r}{\delta t} \) \( = \mathbf F \cdot \mathbf v \)